OpenStudy (clamin):

PLEASE HELP!!! Factor the expression below: 125x^3 + 512y^3

OpenStudy (anonymous):

\(5^3=125,8^3=512\) making this one the "sum of two cubes"

RhondaSommer (rhondasommer):

i can explain it to you ina sec

RhondaSommer (rhondasommer):

can you wait a sec?

OpenStudy (clamin):

@RhondaSommer yes

RhondaSommer (rhondasommer):

here

RhondaSommer (rhondasommer):

alright This fits the form \[a^3+b^3\]

OpenStudy (xapproachesinfinity):

\[a^3+b^3=(a+b)(a^2-ab+b^2)\]

OpenStudy (xapproachesinfinity):

rewrite it that way

RhondaSommer (rhondasommer):

so rewrite it where a=5x and b=8y

RhondaSommer (rhondasommer):

@xapproachesinfinity i got this if you dont mind...

OpenStudy (xapproachesinfinity):

:) i was willing to just you found me doing that

RhondaSommer (rhondasommer):

lol :)

RhondaSommer (rhondasommer):

@clamin rewrite it as \[a^3+b^3\] please

RhondaSommer (rhondasommer):

@clamin you there?

OpenStudy (clamin):

(5x + 8y)

OpenStudy (clamin):

is it this like this?? (5x+8y)(25x^2 - 40xy+ 64y^2)

OpenStudy (anonymous):

\[\large \color\red{a^3}+\color\green{b^3}\]

OpenStudy (anonymous):

looks good to me

OpenStudy (clamin):

is all right??

RhondaSommer (rhondasommer):

\[(5x+8y)(25x^2−40xy+64y^2)\]

RhondaSommer (rhondasommer):

yep looks good

OpenStudy (anonymous):

yes, your answer is correct

OpenStudy (xapproachesinfinity):

great! :)